Prev Next

SW&IT Meeting Minutes 2020-07-20

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20200713 - Bec, Matthieu 40.0
LR20200713 - Holmberg, Carl 38.0
LR20200713 - Edmeades, David 40.0
LR20200713 - Sola, Igor 40.0
LR20200713 - Hooper, Stephen 80.0
LR20200713 - Bryant, Trey 0.0
LR20200713 - Zhang, Yang 40.0
LR20200713 - Zhang, Xianyu 40.0

News from management

From David Carroll

Please be advised that the Campus Health SAFER system to report positive COVID-19 cases is now available here

Note that the SAFER page was referenced on the HR page sent earlier this week under section "Supervisor Required Actions Regarding Employee" (a)(2) ( 

HR update of their protocol page pending.

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • AO data (per last week emails)
  • SSD

Team reports


  • Previous:
    • TMS TCP measurement tests with Heejoo (Monday & Wednesday & Thursday), still having lots of issues.
    • Some other TMS development (wiki, meeting, small code development, etc.).
    • AO-Cravat work on recent OSCO data analysis, plus meeting on Friday.
  • Next:
    • More TMS test with eTalon German Engineers for troubleshooting (tentative schedule on early morning Tuesday).
    • Some more TMS work (code, Ciddor correction study, etc.).
    • AO-Cravat work (OSCO data analysis, sharing results with Julian, code enhancement, etc.).
    • Patching DMS LBC code.
    • IT#8140 (TCS/Adsec slew protection) if have time.
    • Safety training.


  • Last week
    • Log move script mods: direct rsync, exit code checking
    • Laura WFH setup, in good state
      • Asked her to drop off old laptop in Steward next she goes
    • Synology drive swap/replication error stop (?)
      • Broke folder replication because pw change, fixed and whitelisted mountain syno on Tucson. Unfortunately there does not appear to be an official option for SSH keys.
    • Mac mini software installs/dropoff in Steward
    • Testing new VPN/largely great, minor local device access issue noted in Slack
    • Veeam SMTP server now works after SH fixed DNS for vmmgr
  • This week
    • Vacation


  • Last Week:
    • AHS: LBTI alarm patch
    • SupportScripts: replaced h5csv with dms api call in aos_count, less invasively.
    • IT8137: Assisted JPH w/ AGW1 troubleshooting.
  • This Week:
    • AHS: Finish LBTI alarm patch
    • IT8137: Create oacontrol patch with more informative error messages.


Last week:
  • LUCI
    • After uploading new masks in a cabinet exchange, somehow the old masks are still reported. I managed to replicate the problem with a lot of back and forth, trying to replicate the sequence as in the logs. This issue has been there since the beginning. I started developing a solution to it, trying to make it as small as possible.
  • Safety Training

This week:
  • Vacation


  • Last week
    • OC, Management
    • PE's start
    • Recruitment, policy working group
    • SHARK-VIS RFA complete
    • iLocater, AGW, TMS, AO-cravat meetings
    • python-iif/irc hello-world git cleanup
  • This week
    • PE's complete
    • Recruitment, review, questions and phone screens
    • PEPSI new servers follow-up
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png preview
      • max_bytes for ovms


  • Last week
    • Setting up a Github repo for easy access/development
    • Able to now select a 24 hour window and toggle between hour-long 'chunks'
    • 3 separate sets of axes; matplotlib exploration (title/labels overlap still)
    • More efficent querying of DMS data (querying only once instead of twice)
  • This week
    • Still to look at Jupyter Notebook
    • Look at fixes for window scaling w/o keystroke lag
    • RMS calculation; plotting proper data on the 3 sets of axes
    • Fixing overlap of titles/axis labels between the plots
    • Converting datetime timestamps from local time to UTC time


  • VPN account planning, more needed
  • prep for SSD work -- VPN, mountain firewall issues, etc.
  • Steward power last weekend, mostly Ok
    • Synology share permission issues on "/volume1/homes" -- shell, abell and oingo
    • seems to have firewall problem with Radius
    • hv02 is asking for F1 to be pressed before booting.
    • DNS is too entangled on the ovirt cluster.
    • Manually start everything on vm2 (?), also test VMs need to be removed -- IT should keep this clean.
  • CentOS 8 planning -- no go
  • Cisco licensing for ASA.
  • Docker host to ovirt
  • ASC North mark
  • Performance review.


  • This week
    • Safety training
    • SSD support
    • Power Backplane action items
Topic revision: r14 - 23 Jul 2020, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback